OpenStudy (anonymous):

the slope of the line 2x+3y=-9 is?

OpenStudy (anonymous):

-2/3

OpenStudy (anonymous):

present your original equation as: 3y=-2x-9 y=-2/3 *x -3 your slope = number in a front of x... which is = -2/3

OpenStudy (anonymous):

thank you guys!